# YITP-25-175, RESCEU-24/25, IPMU25-0051

### v1

@@ 1 Introduction
Black holes represent one of the most striking predictions of general relativity. In recent years, extremal black holes received considerable attention as they present some qualitatively new phenomena with respect to their non-extremal counterpart [1, 2, 3]. In particular, the stability and regularity of extremal horizons has been analyzed in several different contexts. To start, linear perturbations grow along an extremal horizon in stark contrast with the decay exhibited by perturbation of non-extremal spacetimes [4, 5, 6]. More recently, it was argued that an extremal Reissner–Nordström Anti-de Sitter (AdS) black hole has a singular behavior once perturbed with a scalar field [7]. Subsequently, studying how an extremal Kerr black holes responds to perturbations, it was shown that the properties of extremal horizons can be very sensitive to ultraviolet physics [8] (see also [9]). This effect is even stronger for charged rotating black holes[10]. Finally, similar conclusions hold for extremal black holes in higher dimensions [11].
@@
We start with an analysis of the properties of a test scalar field on the background of general spherically symmetric black-hole spacetimes. Both the extremal and non-extremal cases are studied for the sake of comparison, and we also analyze the case of a dynamical (i.e. time dependent) scalar field. This is done both in static coordinates and in Eddington-Finkelstein-like ones. We then treat, as a very simple particular case, the Reissner–Nordström black holes with a negative cosmological constant and we review some of their properties and some claims of Ref. [7].
@@
Sec. 5:
We end the paper with a brief discussion and an Appendix where some long calculations are summarized.
@@
    Notice that $f_{1}:=f^{\prime\prime}(r_{+})/2$ in this case. Therefore, the $\omega_{\ell m}\to 0$ limit of the index $\gamma$ does not agree with $\gamma$ at $\omega_{\ell m}=0$ ($\lim_{\omega_{\ell m}\to 0}\gamma\neq\gamma|_{\omega_{\ell m}=0}$), showing a kind of discontinuity. Actually, for non-zero but small $\omega_{\ell m}$, the expansion coefficients in (9) behave as $\alpha_{n}/\alpha_{0}=\mathcal{O}(\omega_{\ell m}^{-n})$ and the regime of validity of the series expansion (9) becomes narrower and narrower as $|\omega_{\ell m}|$ decreases, signaling the discontinuity at $\omega_{\ell m}=0$. In general $\gamma^{\ell}_{-}$ is negative and thus must be discarded. Concerning $\gamma^{\ell}_{+}$, for $\ell$ large enough, $\gamma^{\ell}_{+}$ is always greater than or equal to $1$ so that $R_{\ell m}$ and its first $r$-derivative are finite at $r=r_{+}$. For small values of $\ell$, whether $\gamma^{\ell}_{+}\geq 1$ or not depends on $f_{1}r^{2}_{+}$ and thus on the particular extremal black hole. The example provided in [7] requires charge and a negative cosmological constant, so that as we discuss below $\gamma^{1}_{+}<1$, but there may be many different cases.
@@ 2.1 Extremal Reissner–Nordström AdS black hole
The particular case studied in [7] is that of a extremal Reissner–Nordström AdS black hole spacetime. The background geometry is given by (1) with $\beta=0$ and
@@
We now add a test massless scalar field on this background and study static configurations. The above analysis shows that in the near-horizon limit the scalar field behaves as [7]
@@
where $C_{\ell}$ is the amplitude of the field and the value of $\gamma^{\ell}$ is derived from the general relation (13). The exponents $\gamma^{\ell}$ are positive for every choice of the parameters, implying that the scalar field vanishes at the extremal horizon. However, for $\ell=1$ we obtain $0<\gamma^{1}<1$. Therefore, in this case the first $r$-derivative of the scalar field diverges at the horizon. Ref. [7] argues that this leads to a divergent backreaction on the metric. Indeed, we can compute the stress energy tensor for the massless scalar field with the profile of Eq. (26). In general we have products of modes with identical or different $\ell$’s. Nonetheless, since we wish to discuss the most singular (or the least regular) situation, we restrict our considerations to the product of the $\ell=1$ mode with itself. The non-zero components of the stress energy tensor are (with $\gamma$ as a shorthand for $\gamma^{1}$)
@@
Nonetheless, in the natural orthonormal tetrad, none of the components of $T_{\mu\nu}$ diverge. In particular $T_{11}=f(r)T_{rr}$ actually vanishes at the horizon. In fact, all the invariants that can be built from the stress energy tensor are instead finite as the $T_{rr}$ component would necessarily be contracted with the $g^{rr}$ component of the metric which vanishes at the horizon. However, Ref. [7] claims that the divergence of $T_{rr}$ is problematic regardless the finiteness of the invariant quantities. There are at least two reasons why this could be the case according to [7]. First, the divergence of this component would imply a divergent source of backreaction once this form of the stress energy tensor is inserted into the Einstein field equations, leading to a singular spacetime. Second, by using the advanced coordinates $\{v,\bar{r}\}$ of (15) so that the horizon is regularly described, the component
@@ 3.1 Computation of the Backreaction
For simplicity, we only study the case $m=0$. As we are interested in a scalar field as a source, we can restrict our attention to the even sector for the metric perturbation. Furthermore, we are mainly concerned about non-spherical perturbation sourced by the product of the $\ell=1$ scalar field mode with itself. Therefore, we could restrict our consideration to the $\ell=2$ even sector for the metric perturbation. However, we consider a generic $\ell$ ($\geq 2$) mode as that does not increase the degree of complexity. We consider the perturbed metric [12, 13]
@@
Another useful lesson we can learn on the form of the metric after the backreaction without solving the system of equations can be drawn by looking at the $\{r,\theta\}$ equation and noticing that it is an homogeneous equation in the perturbation $h^{\ell}(r)$. Therefore, there exist solutions for which $g_{r\theta}$ vanishes and the metric remains diagonal. For simplicity, we focus on this class of geometries in the rest of the work. Furthermore, it is always possible to find a coordinate system in which $G^{\ell}(r)=0$ (see e.g., [14]). In the rest of this work, we consider a geometry compatible with the expressions in (33) with $G^{\ell}(r)=h^{\ell}(r)=0$, but which is not necessarily obtained as a result of computation performed in the realm of perturbation theory.
@@ 3.2 Geodesic completeness
However, the right hand side of Eq. (41) generically depend on $\theta$, thus it does not correspond to a well defined coordinate transformation. Furthermore, even if we restrict to the case where the ratio on the right hand side of Eq. (41) is $\theta$ independent, if we now perform a coordinate transformation to express the metric in terms of the coordinate (40), we immediately see that we do not obtain a set of coordinates regular at the horizon. This explains that the divergence of the Raychaudhuri equation noted in [7] is not necessarily due to the presence of a physical singularity, but it might simply indicate that we have yet to identify a proper suitable coordinate set to describe the horizon.
@@
There is the typical divergence at the horizon for $\dot{t}$, at finite affine parameter, implying that $t$ is not a good coordinate. However, there is a more worrying problem at the equation for $\ddot{\theta}$, because the last term on the righthand side of Eq. (64) diverges at the horizon, unless $FH=R^{2}(r)$, and this cannot be cured with any extension of the coordinate $\theta$. Therefore, we conclude that, despite the absence of scalar curvature singularities, spacetimes that do not satisfy Eq. (56) are pathological as there are incomplete geodesics.
@@ 4.1 Energy conditions
which can be brought to a typical form of the Plebanski-Demianski general family of solutions (see e.g. the recent [15, 16]). By performing the coordinate change
@@
It is well known [17] that this horizon is regular except for a deficit angle at one of the poles $\theta=0,\pi$ —this deficit angle is present in the whole static external region. This is a rather mild problem and has nothing to do with the kind of divergences predicted in [7]. For the general metric (65) the avoidance of any deficit angles at both poles requires that the following condition
@@
In our case, we wish to consider the extremal cases. For the metric (66) with (70) they arise when the function $R(r)$ has a double positive root. To know the number of positive roots we use Descartes’s rule of signs, so that there are either 3 or 1 positive roots if $\Lambda+3\alpha^{2}>0$, and there are 2 or none positive roots if $\Lambda+3\alpha^{2}\leq 0$. In all cases, when two of the positive roots coincide one has the extremal case, see [18]. The marginal case with $\Lambda+3\alpha^{2}=0$ is particularly interesting because, on one hand this solution with $Q=0$ has been used to describe a single black hole on the brane [19], and on the other hand $R(r)=0$ becomes a cubic equation and the conditions for a double root are easily found. The restrictions on the parameters for the existence of the double root in this marginal case are
@@ 4.2 Scalar field in the extremal RN-AdS black hole
Once we have proven that there are perturbations of extremal BHs of type (36) that keep an extremal regular horizon and satisfy the energy conditions, we would like to know if such a perturbation can be due to the presence of a massless minimally coupled scalar field. To that end we exploit the result of corollary 6.3 in Ref. [20] which provides the algebraic conditions that the energy-momentum tensor corresponding to a minimally coupled scalar field must have. In particular, according to the corollary, a symmetric tensor $T_{ab}$ is algebraically the energy-momentum tensor of a minimally coupled massless scalar field if and only if
@@
and we can apply the Corollary 6.3 of [20] to the lefthand side of this relation. This is done in the Appendix, where we found it much easier to perform the analysis in the coordinate system of the metric form (69) and where, as a consistency check, we controlled the calculations keeping in mind that there is a solution with all $g_{n}=h_{n}=0$ for $n\geq 1$ and with
@@
The long but straightforward computation summarized in the Appendix shows that the constraint of Eq. (72) imposed by the Corollary 6.3 of [20] can be, in the near horizon limit, solved by the stress energy tensor in Eq. (82). In other words, the equalities in Eq. (72) do hold close to an extremal horizon assuming that (79) describes the electromagnetic field for the perturbed solution as well. This shows that, under this assumption, black holes with regular deformed horizons can be produced by perturbing a Reissner–Nordström AdS black hole with a scalar field —at least near the horizon.
@@ 5 Conclusions
Motivated by the qualitative new features that arise when considering extremal black holes compared to their non-extremal counterpart, we have analyzed the regularity of extremal Reissner–Nordström AdS black holes under scalar perturbations. As it was previously noted [7], the static solution of scalar wave equation leads to a stress energy tensor with some diverging component. We have shown that this divergence does not lead to a divergent backreaction as the potentially problematic divergent behavior of $T_{rr}$ is compensated by the vanishing behavior of the background metric $g_{rr}$ at the horizon. In fact, all curvature scalars remain finite and the backreaction of the scalar field on the geometry can be studied within the framework of perturbation theory.
@@
The crucial remaining question concerns the existence of physically reasonable sources for these constrained geometries. We showed that spacetimes satisfying the above regularity condition include well-known solutions in the Plebanski–Demiański class, such as the charged C-metric with cosmological constant, which obey standard energy conditions. Finally, we investigate whether the perturbation due to the scalar field of an extremal Reisner–Nordström black hole leads to a well defined horizon that satisfy the regularity constraint. While we were not able to fully answer this question due to the fact that we cannot consistently solve the backreaction equations everywhere, we showed that, at least locally, there exists solutions of deformed extremal horizons due to the presence of the scalar field that are perfectly regular. As the analysis only focuses on the near horizon region, there is no guarantee that it is possible to find an everywhere well-defined static geometry. If we were able to find an everywhere defined solution that describes deformed extremal black holes and that is asymptotically flat (or asymptotically AdS), the tidal Love numbers of the undeformed black hole [21, 22] would be infinite. While this is technically possible, this type of solution is definitely non-standard and there are no similar solutions in the literature. Therefore, it is maybe more likely that the class of extremal horizon geometry that is locally regular corresponds to a new class of spacetimes representing extremal black holes deformed by external gravitational sources. Depending on the timescale of physical interest, the region between the deformed extremal horizon and the external gravitational source may be approximated by the static geometry discussed here. On the other hand, the regions near and outside the external gravitational sources are probably not.
Overall, our results show that extremal horizons in AdS are not generically singular under scalar perturbations. Instead, there exists a nontrivial family of deformed extremal characterized by a specific geometric constraint. Furthermore, the violation of this constraint leads to the presence of incomplete geodescs at the horizon, not divergence of curvature invariants. Therefore, extremal black holes do not inherently act as universal amplifiers of new physics.

### v2

@@ 1 Introduction
Black holes represent one of the most striking predictions of general relativity. In recent years, extremal black holes received considerable attention as they present some qualitatively new phenomena with respect to their non-extremal counterpart [Hawking:1994ii, Kunduri:2013gce, Dafermos:2025int]. In particular, the stability and regularity of extremal horizons has been analyzed in several different contexts. To start, linear perturbations grow along an extremal horizon in stark contrast with the decay exhibited by perturbation of non-extremal spacetimes [Aretakis:2011ha, Aretakis:2011gz, Aretakis:2012ei]. More recently, it was argued that extremal Reissner–Nordström Anti-de Sitter (AdS) black holes have a singular behavior once perturbed with a scalar field [Horowitz:2022mly]. Similarly, studying how an extremal Kerr black holes responds to perturbations and quantum corrections of charged black holes, it was shown that the properties of extremal horizons can be very sensitive to ultraviolet physics [Horowitz:2023xyl, Cano:2024bhh, Chen:2024sgx] (see also [DelPorro:2025fiu]). This effect is even stronger for charged rotating black holes[Horowitz:2024dch]. Finally, similar conclusions hold for extremal black holes in higher dimensions [Horowitz:2024kcx].
@@
We start with an analysis of the properties of a test scalar field on the background of general spherically symmetric black-hole spacetimes. Both the extremal and non-extremal cases are studied for the sake of comparison, and we also analyze the case of a dynamical (i.e. time dependent) scalar field. This is done both in static coordinates and in Eddington-Finkelstein-like ones. We then treat, as a very simple particular case, the Reissner–Nordström black holes with a negative cosmological constant and we review some of their properties and some claims of Ref. [Horowitz:2022mly].
@@
Conclusions:
We end the paper with a brief discussion of the main result and two appendixes. In App A some long calculations are summarized, while in App. B we briefly discuss the higher dimensional case.
@@ Conclusions:
    Notice that $f_{1}:=f^{\prime\prime}(r_{+})/2$ in this case. Therefore, the $\omega_{\ell m}\to 0$ limit of the index $\gamma$ does not agree with $\gamma$ at $\omega_{\ell m}=0$ ($\lim_{\omega_{\ell m}\to 0}\gamma\neq\gamma|_{\omega_{\ell m}=0}$), showing a kind of discontinuity. Actually, for non-zero but small $\omega_{\ell m}$, the expansion coefficients in (9) behave as $\alpha_{n}/\alpha_{0}=\mathcal{O}(\omega_{\ell m}^{-n})$ and the regime of validity of the series expansion (9) becomes narrower and narrower as $|\omega_{\ell m}|$ decreases, signaling the discontinuity at $\omega_{\ell m}=0$. In general $\gamma^{\ell}_{-}$ is negative and thus must be discarded. Concerning $\gamma^{\ell}_{+}$, for $\ell$ large enough, $\gamma^{\ell}_{+}$ is always greater than or equal to $1$ so that $R_{\ell m}$ and its first $r$-derivative are finite at $r=r_{+}$. For small values of $\ell$, whether $\gamma^{\ell}_{+}\geq 1$ or not depends on $f_{1}r^{2}_{+}$ and thus on the particular extremal black hole. The example provided in [Horowitz:2022mly] requires charge and a negative cosmological constant, so that as we discuss below $\gamma^{1}_{+}<1$, but there may be many different cases. [^1] The analysis can be generalized to higher dimensions as shown in App. B (see also [Chen:2024sgx, Acharya:2025ojn]).
@@ 2.1 Extremal Reissner–Nordström AdS black hole
The particular case studied in [Horowitz:2022mly] is that of a extremal Reissner–Nordström AdS black hole spacetime. The background geometry is given by (1) with $\beta=0$ and
@@
We now add a test massless scalar field on this background and study static configurations. The above analysis shows that in the near-horizon limit the scalar field behaves as [Horowitz:2022mly]
@@
where $C_{\ell}$ is the amplitude of the field and the value of $\gamma^{\ell}$ is derived from the general relation (13). The exponents $\gamma^{\ell}$ are positive for every choice of the parameters, implying that the scalar field vanishes at the extremal horizon. However, for $\ell=1$ we obtain $0<\gamma^{1}<1$. Therefore, in this case the first $r$-derivative of the scalar field diverges at the horizon. Ref. [Horowitz:2022mly] argues that this leads to a divergent backreaction on the metric. Indeed, we can compute the stress energy tensor for the massless scalar field with the profile of Eq. (26). In general we have products of modes with identical or different $\ell$’s. Nonetheless, since we wish to discuss the most singular (or the least regular) situation, we restrict our considerations to the product of the $\ell=1$ mode with itself. The non-zero components of the stress energy tensor are (with $\gamma$ as a shorthand for $\gamma^{1}$)
@@
Nonetheless, in the natural orthonormal tetrad, none of the components of $T_{\mu\nu}$ diverge. In particular $T_{11}=f(r)T_{rr}$ actually vanishes at the horizon. In fact, all the invariants that can be built from the stress energy tensor are instead finite as the $T_{rr}$ component would necessarily be contracted with the $g^{rr}$ component of the metric which vanishes at the horizon. However, Ref. [Horowitz:2022mly] claims that the divergence of $T_{rr}$ is problematic regardless the finiteness of the invariant quantities. There are at least two reasons why this could be the case according to [Horowitz:2022mly]. First, the divergence of this component would imply a divergent source of backreaction once this form of the stress energy tensor is inserted into the Einstein field equations, leading to a singular spacetime. Second, by using the advanced coordinates $\{v,\bar{r}\}$ of (15) so that the horizon is regularly described, the component
@@ 3.1 Computation of the Backreaction
For simplicity, we only study the case $m=0$. As we are interested in a scalar field as a source, we can restrict our attention to the even sector for the metric perturbation. Furthermore, we are mainly concerned about non-spherical perturbation sourced by the product of the $\ell=1$ scalar field mode with itself. Therefore, we could restrict our consideration to the $\ell=2$ even sector for the metric perturbation. However, we consider a generic $\ell$ ($\geq 2$) mode as that does not increase the degree of complexity. We consider the perturbed metric [Regge:1957td, PhysRevLett.24.737]
@@
Another useful lesson we can learn on the form of the metric after the backreaction without solving the system of equations can be drawn by looking at the $\{r,\theta\}$ equation and noticing that it is an homogeneous equation in the perturbation $h^{\ell}(r)$. Therefore, there exist solutions for which $g_{r\theta}$ vanishes and the metric remains diagonal. For simplicity, we focus on this class of geometries in the rest of the work. Furthermore, it is always possible to find a coordinate system in which $G^{\ell}(r)=0$ (see e.g., [DeFelice:2024dbj]). In the rest of this work, we consider a geometry compatible with the expressions in (33) with $G^{\ell}(r)=h^{\ell}(r)=0$, but which is not necessarily obtained as a result of computation performed in the realm of perturbation theory.
@@ 3.2 Geodesic completeness
However, the right hand side of Eq. (41) generically depend on $\theta$, thus it does not correspond to a well defined coordinate transformation. Furthermore, even if we restrict to the case where the ratio on the right hand side of Eq. (41) is $\theta$ independent, if we now perform a coordinate transformation to express the metric in terms of the coordinate (40), we immediately see that we do not obtain a set of coordinates regular at the horizon. This explains that the divergence of the Raychaudhuri equation noted in [Horowitz:2022mly] is not necessarily due to the presence of a physical singularity, but it might simply indicate that we have yet to identify a proper suitable coordinate set to describe the horizon.
@@
There is the typical divergence at the horizon for $\dot{t}$, at finite affine parameter, implying that $t$ is not a good coordinate. However, there is a more worrying problem at the equation for $\ddot{\theta}$, because the last term on the righthand side of Eq. (64) diverges at the horizon, unless $FH=R^{2}(r)$, and this cannot be cured with any extension of the coordinate $\theta$. Therefore, we conclude that, despite the absence of scalar curvature singularities, spacetimes that do not satisfy Eq. (56) are pathological as there are incomplete geodesics. This condition is also necessary for higher dimensional spacetimes as shown in Appendix B
@@ 4.1 Energy conditions
which can be brought to a typical form of the Plebanski-Demianski general family of solutions (see e.g. the recent [Ovcharenko:2024yyu, Ovcharenko:2025fxg]). By performing the coordinate change
@@
It is well known [Griffiths:2009dfa] that this horizon is regular except for a deficit angle at one of the poles $\theta=0,\pi$ —this deficit angle is present in the whole static external region. This is a rather mild problem and has nothing to do with the kind of divergences predicted in [Horowitz:2022mly]. For the general metric (65) the avoidance of any deficit angles at both poles requires that the following condition
@@
In our case, we wish to consider the extremal cases. For the metric (66) with (70) they arise when the function $R(r)$ has a double positive root. To know the number of positive roots we use Descartes’s rule of signs, so that there are either 3 or 1 positive roots if $\Lambda+3\alpha^{2}>0$, and there are 2 or none positive roots if $\Lambda+3\alpha^{2}\leq 0$. In all cases, when two of the positive roots coincide one has the extremal case, see [Dias:2002mi]. The marginal case with $\Lambda+3\alpha^{2}=0$ is particularly interesting because, on one hand this solution with $Q=0$ has been used to describe a single black hole on the brane [Emparan:1999wa], and on the other hand $R(r)=0$ becomes a cubic equation and the conditions for a double root are easily found. The restrictions on the parameters for the existence of the double root in this marginal case are
@@ 4.2 Scalar field in the extremal RN-AdS black hole
Once we have proven that there are perturbations of extremal BHs of type (36) that keep an extremal regular horizon and satisfy the energy conditions, we would like to know if such a perturbation can be due to the presence of a massless minimally coupled scalar field. To that end we exploit the result of corollary 6.3 in Ref. [Bergqvist:2001rz] which provides the algebraic conditions that the energy-momentum tensor corresponding to a minimally coupled scalar field must have. In particular, according to the corollary, a symmetric tensor $T_{ab}$ is algebraically the energy-momentum tensor of a minimally coupled massless scalar field if and only if
@@
and we can apply the Corollary 6.3 of [Bergqvist:2001rz] to the lefthand side of this relation. This is done in the Appendix, where we found it much easier to perform the analysis in the coordinate system of the metric form (69) and where, as a consistency check, we controlled the calculations keeping in mind that there is a solution with all $g_{n}=h_{n}=0$ for $n\geq 1$ and with
@@
The long but straightforward computation summarized in the Appendix shows that the constraint of Eq. (72) imposed by the Corollary 6.3 of [Bergqvist:2001rz] can be, in the near horizon limit, solved by the stress energy tensor in Eq. (82). In other words, the equalities in Eq. (72) do hold close to an extremal horizon assuming that (79) describes the electromagnetic field for the perturbed solution as well. This shows that, under this assumption, black holes with regular deformed horizons can be produced by perturbing a Reissner–Nordström AdS black hole with a scalar field —at least near the horizon.
@@ 5 Conclusions
Motivated by the qualitative new features that arise when considering extremal black holes compared to their non-extremal counterpart, we have analyzed the regularity of extremal Reissner–Nordström AdS black holes under scalar perturbations. As it was previously noted [Horowitz:2022mly], the static solution of scalar wave equation leads to a stress energy tensor with some diverging component. We have shown that this divergence does not lead to a divergent backreaction as the potentially problematic divergent behavior of $T_{rr}$ is compensated by the vanishing behavior of the background metric $g_{rr}$ at the horizon. In fact, all curvature scalars remain finite and the backreaction of the scalar field on the geometry can be studied within the framework of perturbation theory.
@@
The crucial remaining question concerns the existence of physically reasonable sources for these constrained geometries. We showed that spacetimes satisfying the above regularity condition include well-known solutions in the Plebanski–Demiański class, such as the charged C-metric with cosmological constant, which obey standard energy conditions. Finally, we investigate whether the perturbation due to the scalar field of an extremal Reisner–Nordström black hole leads to a well defined horizon that satisfy the regularity constraint. While we were not able to fully answer this question due to the fact that we cannot consistently solve the backreaction equations everywhere, we showed that, at least locally, there exists solutions of deformed extremal horizons due to the presence of the scalar field that are perfectly regular. As the analysis only focuses on the near horizon region, there is no guarantee that it is possible to find an everywhere well-defined static geometry. If we were able to find an everywhere defined solution that describes deformed extremal black holes and that is asymptotically flat (or asymptotically AdS), the tidal Love numbers of the undeformed black hole [Damour:2009vw, Binnington:2009bb] would be infinite. While this is technically possible, this type of solution is definitely non-standard and there are no similar solutions in the literature. Therefore, it is maybe more likely that the class of extremal horizon geometry that is locally regular corresponds to a new class of spacetimes representing extremal black holes deformed by external gravitational sources. Depending on the timescale of physical interest, the region between the deformed extremal horizon and the external gravitational source may be approximated by the static geometry discussed here. On the other hand, the regions near and outside the external gravitational sources are probably not.
Overall, our results show that extremal horizons in AdS are not necessarily singular under scalar perturbations. Instead, there exists a nontrivial family of deformed extremal characterized by a specific geometric constraint. Furthermore, the violation of this constraint leads to the presence of incomplete geodescs at the horizon, not divergence of curvature invariants. Therefore, this class of extremal black holes do not inherently act as universal amplifiers of new physics.